OpenStudy (anonymous):

2/5(x - 4) = 2x for x. Answer 1 -1/2 -1 -8

OpenStudy (anonymous):

it is -1 if u just need answer

OpenStudy (mr.math):

Multiply both sides by 5 for simplification: \[2(x-4)=10x\] Now open the parenthesis and add the x's together: \[2x-8=10x \implies 8x=-8 \implies x=-1\]

OpenStudy (anonymous):

2/5x-8/5=2x x=1/5x-4/5 4/5x=-4/5 x=-1